# Detection of the papermilling behavior

### v2

@@ Abstract
1 Introduction: the good intentions
The original idea of publications and citations indexes was to help scientists in their research work. Eugene Garfield, the creator of the Science Citation Index, wrote:
“In this paper, I propose a bibliographic system for science literature that can eliminate the uncritical citation of fraudulent, incomplete, or obsolete data by making it possible for the conscientious scholar to be aware of criticisms of earlier papers.” [1].
However, there is a proverb that says good intentions pave the road to hell, and the field of mathematics is currently an identified victim of massive scientific misconduct related to manipulations with publications and citations counts considered as indicators of the quality of individual researchers. Clearly, other fields of science suffer similar problems, although perhaps less visible.
2 The current situation: “code red”
In the description of the methodology for preparing the Highly Cited Researchers™  2023 lists, Clarivate™ explains the reasons for excluding the field of mathematics from such list [2, see section “Exceptions and exclusions”]:
@@ 2 The current situation: “code red”
It should be added that the lists of highly cited researchers in mathematics contained 99 names in 2014, also 99 names in 2015, 106 names in 2016, 96 names in 2017, 90 names in 2018, 89 names in 2019, 70 names in 2020, 74 names in 2021, 52 names in 2022, and no names after that [3]. These lists are based on the data from the preceding years – for example, the 2022 list is based on the data from 2011 to 2021. In all the lists there are 243 unique names.
@@
It is worth mentioning here, that the fact that publications and citations counts in mathematics are the lowest compared to other sciences is known, and has been used for comparing citation impact in different fields of science [4] and for creating the first multidisciplinary list of highly cited researchers based on their citation counts normalized with respect to mathematics [5].
@@
At the time of writing of the paper [5], the Highly Cited Researchers list (at the time known as Thomson ISI) was compiled based on the citations counts over the duration of the entire career. There was also a category called “Research front”, where were included papers that collected some threshold number of citations within short time (a couple of years only); indeed, this could indicate that such papers might be bringing really new or significant results.
@@
Currently, Clarivate compiles its Highly Cited Researchers (HCR) lists using something similar to those previous “Research front” papers [2, see section “Preliminary candidate selection”]:
@@
“For the Highly Cited Researchers 2023 analysis, the papers surveyed were those published during 2012 to 2022 and which then ranked in the top 1% by citations for their ESI field and year (the definition of a highly cited paper).”
The abbreviation ESI means Essential Science Indicators™  – another Clarivate product.
@@
Instead of the accumulated whole-career citation impact, the “beans” called “highly cited papers” are counted for a relatively short time window. Due to this change of rules, it became possible – exactly as Clarivate noticed – to arrange relatively small numbers of citations for a set of several papers and, as a result, get a particular person to the HCR list. The field of mathematics appeared particularly vulnerable to this type of action, because of its low citations counts per paper.
@@
However, mathematics has the intrinsic property of a usually long delay between the publication of a result and its wide recognition and use, reflected in the number of citations. To artificially increase the number of citations for the produced/selected “beans” within a short time frame, the manipulation of citations is used used, which is done by producing more insignificant papers with the sole purpose of increasing the citation counts for the “beans”.
@@
Apparently, the people who become the “recipients” of a papermill in order to increase their publication counts, also become the “providers” of citations for other recipients through their subsequent papermill publications. This leads to the snowball effect, and is likely the reason for the rapid monotonous growth in publication counts of individuals involved in the papermilling process.
@@ 3 The data source
The data for further processing are obtained by exporting necessary search results from the Web of Science (WoS). The necessary steps are as follows.
@@
4 Visual signs of papermilling behavior
@@
This plot allows for hovering over or clicking on the vertical bars (or on the line) in order to see the number of publications and citations in a particular year.
Two sample outputs are shown in Fig. 1 (researcher R1) and Fig. 2 (researcher R2), and in both cases the researchers are mathematicians who have been listed in the Highly Cited Researchers lists in the past; their career lengths are similar.
@@
However, from 2011 on, the number of publications of R2 started rapidly growing, giving the total of 303 publications in 2011–2023, reaching the peak of 45 papers in 2022, and the number of citations started to grow synchronously with the number of publications. The total number of co-authors of R2 during 2011–2023 was 148, from 35 countries.
The described differences between researchers R1 and R2 are obvious. It should be mentioned that the publication pattern of R1 also reflects the generally-known observation that significant results are achieved in the first half of a mathematician’s career, and that the publication activity normally decreases with age. In the case of the researcher R2, one can observe the aforementioned snowball effect typical for the papermilling behavior.
Obviously, in the case of researcher R2, authoring 45 papers in mathematics in one year is not typical, and the synchronous growth of citation counts alongside the growth in number of papers is also atypical. On the contrary, in the case of researcher R1, the continuous growth of citations counts despite the decrease in the number of papers is normal and shows the long-lasting impact of R1’s publications in the first half of the considered time period.
The following features in the plots for publications and citations counts are possible indicators of the papermilling behavior:
@@
  • starting from some particular year, the number of publications is starts rapidly growing (usually monotonically), and reaches a number of papers per year which is unrealistic for a conscientious and responsible scientist;
@@
Besides these symptomatic signs of the overall publishing behavior, there are also symptoms that symptoms that require scrutiny of each particular publication.
@@ 5 Basic quantitative indicators of the papermilling behavior
The number of publications for each year and the number of citations in the same year are two time series, and we can investigate their properties and mutual relationship. To quantify the previous observations, the following indicators can be used when examining the publication activity of researchers:
  • correlation between the publications counts and the citations counts; high correlation coefficient (above 0.7–0.8) indicates possible papermilling;
  • the delay between the citations counts and the publications counts; if the correlation is high and the delay is small (less than 2 years), then this indicates possible papermilling;
  • large maximum number of publications in one year; unusually large number of publications in one year, with increasing trend, indicates possible papermilling;
  • low integrity index (see the next Section 6).
@@ 6 Integrity index
The main goal of a conscientious scientist is to discover new knowledge, and not to increase the number of his/her publications. In this respect, it is worth recalling Theodore von Kármán’s approach to comparing scientists [6, page 4]:
“If you define a great scientist as a man with great ideas, then you will have to rate Einstein first – he had four great ideas. In the history of science perhaps only Sir Isaac Newton is ahead of Einstein, because he had five or six ideas.
All the other major scientists of our age are associated with just one, or at the most two great ideas. In my case I have had three great ideas. Maybe more. Yes, perhaps three and a half great ideas.”
The growing numbers of published papers make increasingly difficult to distinguish the valuable, novel, and original material from material that just increases the number of published papers.
The motivation for introducing the h-index by J. E. Hirsch [7] was to focus on the publications which received sufficient attention of the related scientific community reflected in citations, and not to focus on large numbers of publications that did not draw notable attention. However, the papermills allow a fast increase of the h-index by generating a large number of worthless papers and citations to such papers, also in worthless articles; this is the aforementioned snowball effect.
The solution can be introducing a quantitative indicator, which would implement von Kármán’s idea by lowering the motivation for increasing the number of publications by penalizing the unnecessary increase of the number of publications.
As such indicator, I suggest the ratio of the h-index and the total number of publications. This can be considered as a kind of a signal-to-noise ratio, and can be called the integrity index and denoted as “$I$-index”.
If one author with h-index $H=32$ has 64 papers indexed in the Web of Science and the other author with the same h-index $H=32$ has 320 papers, then the integrity index of the first author is $I_{1}=32/64=0.5$ and the integrity index of the second author is $I_{1}=32/320=0.1$. This means the higher probability that the first author focuses on publishing novel original results and not on just publishing another paper, and the lower probability of papermilling compared to the second author.
In the example discussed in Section 4, the integrity index of researcher R1 is $I_{1}=86/196=0.44$, while the integrity index of researcher R2 is $I_{2}=34/384=0.08$. The integrity index can be used to differentiate conscientious scientists from the possible papermillers.
The values of the $I$-index of conscientious scientists are typically higher than 0.2, and in the case of papermilling the integrity index is very low compared to scientists with the similar values of h-index.
@@ 7 The MATLAB function
A function for MATLAB, “researcherprofilewos.m”, [8] has been created to simplify analysis of the data obtained from the Web of Science. The number of publications for each year and the number of citations for each year are just two times series.
@@
  • maximal number of publications in one year;
@@
  • $I$-index (integrity index);
@@
  • total citations count.
@@
  • minimal number of publications in one year;
  • average number of publications in one year;
  • average number of citations per publication.
@@
The key output is the coefficient of linear correlation between publications and citations. Strong correlation means that the citation dynamics is strongly linked to the publication dynamics. In the case of researcher R2 the coefficient of linear correlation between publications and citations is $r_{2}=0.94$, and there is no delay between the growth of the number of publications and the number of citations. For researcher R1, the correlation coefficient is small negative number $r_{1}=-0.23$, which means that R1’s citations counts grow despite decreasing publications counts.
8 Papermilling behavior for promotions?
@@
There are various reasons for papermilling, and being listed in the lists of highly-cited researchers is only one of them. The other reason is financial, because many universities link salaries and other financial incentives to the number of publications and citations. One of the most widespread reasons is pursuing possible promotion. Usually a candidate aspiring for promotion has to have a certain number of publications and citations in journals and conferences indexed in Web of Science or Scopus databases. The desire to achieve a promotion in a short time encourages some researchers to engage in papermilling, because doing conscientious research and waiting for the eventual future citation impact delays the desired promotion.
@@
Two probable examples of such kind are shown in Fig. 3 (mathematics, Czech Republic), and Fig. 4 (mixed bag of environmental science, business, etc., Slovakia). with high correlation between publications and citations.
@@
The damage done by the promotion-motivated papermilling should not be underestimated. This type of behavior is contagious and may strongly affect colleagues and students, pushing them to similar wrong and unethical practices. In addition, the promotion-motivated papermilling frequently helps such papermillers to get nto influential positions or committees where they make important decisions about hiring, grants, and promotions of honest and conscientious researchers – and also of researchers similar to themselves.
Figure 3: Papermilling for promotion: example 1 (mathematics, CZ).
Figure 4: Papermilling for promotion: example 2 (environmental science, SK).
9 Papermilling behavior outside mathematics
The exclusion of the field of mathematics from the Clarivate’s lists of highly cited researchers does not mean that other fields are spared of papermilling, as the following two examples from Slovakia illustrate.
In Fig. 5 a well-documented example from the field of plant science is shown, with very low performance in 1989–2013 and rapid growth after that: 11 papers in 2017, 14 in 2018, 19 in 2019, 33 in 2020, peaking at 90 papers in 2021, accompanied by the growth of citations counts from roughly 30 citations in 2013 to roughly 3100 citations in 2022. The number of co-authors is 1251, from 56 countries.
@@ 9 Papermilling behavior outside mathematics
In Fig. 6 an example from the field of engineering is shown, with very low performance until 2016, and extremely rapid growth after that: 3 papers and 23 citations in 2017, 11 papers and 56 citations in 2018, 32 papers and 307 citations in 2019, more than 13 papers and around 1300 citations in 2020, more than 110 papers and more than 2500 citations in 2021, 96 papers and more than 4300 citations in 2022. The number of co-authors is 1084, coming from 63 countries.

### v4

@@ Abstract
Based on the analysis of the data obtainable from the Web of Science publication and citation database, typical signs of possible papermilling behavior are described, quantified, and illustrated by examples. A MATLAB function is provided for the analysis of the outputs from the Web of Science. A new quantitative indicator – integrity index, or I-index – is proposed for using it along with standard bibliographic and scientometric indicators. A case study is presented.
1 Introduction
Two recent articles by Agricola et al. [2, 3] draw attention to the problems of scientific misconduct that during recent years affected the mathematical community.
There is one important aspect, however, that they did not mention. All those fraudulent papers are submitted to journals by individuals pursuing not the mathematical discoveries, not the advancement of knowledge – but obtaining financial profit, promotions, grants, positions in editorial boards, various awards, including being named “highly cited researchers” (HCR).
@@ 1 Introduction
Identification of papers generated (or produced) by “paper mills” is an important problem. The description of the operation of “paper mills” can be found in [1] and at websites focusing on various forms of scientific misconduct, e.g. [8, 14, 15].
This paper, however, is devoted to the identification of the systematic papermilling behavior of individual persons. It appears that it is possible to tell conscientious scientists from the individuals whose only goal is the increase of their numbers of publications, citations count, and h-index.
The paper is organized as follows. First, the problem with excluding Mathematics from the lists of highly cited researchers published yearly by Clarivate, is described, and it is explained how mathematics became an easy victim of papermilling. Then the visual signs of the papermilling behavior are described and quantified with the help of the MATLAB toolbox created for processing the publication and citation data downloaded from the Web of Science database. As a case study, the 2019 list of highly cited researchers in the field of mathematics is analyzed.
A new indicator, named the integrity index (I-index), is proposed with the aim to help differentiate conscientious scientists and those who exhibit the papermilling behavior.
Although this study started with the close look at the field of mathematics as a victim of papermilling, the papermilling behavior can be observed in other fields, and examples of that are provided.
Finally, the relationship between the introduction of article processing charges (APC) and the open access fees (OAF), on one side, and the observed massive papermilling behavior, on the other, is discussed.
2 The good intentions…
The original idea of publications and citations indexes was to help scientists in their research work. Eugene Garfield, the creator of the Science Citation Index, wrote [6]:
“In this paper, I propose a bibliographic system for science literature that can eliminate the uncritical citation of fraudulent, incomplete, or obsolete data by making it possible for the conscientious scholar to be aware of criticisms of earlier papers.”.
However, there is a proverb that says that good intentions pave the road to hell, and the field of mathematics is currently an identified victim of massive scientific misconduct related to manipulations with publications and citations counts considered as indicators of the quality of individual researchers. Clearly, other fields of science suffer similar problems.
3 …and what happened
In the description of the methodology for preparing the Highly Cited Researchers™ 2023 lists, Clarivate™ explains the reasons for excluding the field of mathematics from such list [4, see section “Exceptions and exclusions”]:
@@ 2 The good intentions…
It should be added that the lists of highly cited researchers in mathematics contained 99 names in 2014, also 99 names in 2015, 106 names in 2016, 96 names in 2017, 90 names in 2018, 89 names in 2019, 70 names in 2020, 74 names in 2021, 52 names in 2022, and no names after that ([5]). These lists are based on the data from the preceding years – for example, the 2022 list is based on the data from 2011 to 2021. In all the lists there are 243 unique names. However, not all of them have unique Web of Science author identifiers.
@@
It is worth mentioning here, that the fact that publications and citations counts in mathematics are the lowest compared to other sciences is known, and has been used for comparing citation impact in different fields of science ([10, 13]) and for creating the first multidisciplinary list of highly cited researchers based on their citation counts normalized with respect to mathematics ([12]).
@@
At the time of writing of the paper [12], the Highly Cited Researchers list (at the time known as Thomson ISI) was compiled based on the citations counts over the duration of the entire career. There was also a category called “Research front” including those papers that received more than some threshold number of citations within short time (a couple of years only); indeed, this could indicate that such papers might be bringing really new or significant results.
@@
Currently, Clarivate compiles its Highly Cited Researchers (HCR) lists using something similar to those previous “Research front” papers [4, see section “Preliminary candidate selection”]:
@@
“For the Highly Cited Researchers 2023 analysis, the papers surveyed were those published during 2012 to 2022 and which then ranked in the top 1% by citations for their ESI [^1] Essential Science Indicators™ – another Clarivate product. field and year (the definition of a highly cited paper).”
@@
This means that instead of the accumulated whole-career citation impact, the “beans” called “highly cited papers” are counted for a relatively short time window of ten years. Due to this change of rules, it became possible – exactly as Clarivate described – to arrange relatively small numbers of citations for a set of several papers and, as a result, get a particular person to the HCR list. The field of mathematics appeared particularly vulnerable to this type of action, because of its low citations counts per paper. Indeed, mathematics has the intrinsic property of a usually long delay between the publication of a result and its wide recognition and use, reflected in the number of citations.
@@
To artificially increase the number of citations for the produced/selected “beans” within a short time frame, in order to obtain various benefits, the manipulation of citations is used, which is done by producing more insignificant papers with the sole purpose of increasing the citation counts for the chosen “beans”. This can be done easily because of the Clarivate’s imperfect methodology and the way of counting citations.
According to Clarivate’s in table of 1% citation thresholds for mathematics, a paper published in 2025 is considered highly cited if it received 3 (three) citations, a paper published in 2024 needs 9 citations, 2023 paper – 19 citations, 2022 paper – 30 citations, 2021 paper – 42 citations, 2020 paper – 56 citations, 2019 paper – 64 citations, 2018 paper – 74 citations, 2017 paper– 77 citations, 2016 – 81 citations, and a 2015 paper needs 92 citations for being considered as highly cited. When counting these citations, Clarivate does not ignore self-citations by all co-authors, and even if only the co-authors cited their joint paper in three other papers indexed by Clarivate, that joint paper becomes “highly cited”. This methodology and low thresholds allows coordinated action of groups of researchers for producing bunches of papers in co-authorship citing each other.
@@
Apparently, those people who join a papermilling group in order to increase their publication counts, also become the “suppliers” of citations for other members of the group through their subsequent papermilling publications. This leads to the snowball effect, and is likely the reason for the rapid monotonic growth of publication counts of individuals involved in the papermilling process. It also leads to the growth of the number of co-authors of such individuals.
@@ 4 The data source
The data for further processing are obtained by exporting search results from the Web of Science (WoS). The necessary steps are the following.
@@
5 Papermilling behavior: visual signs
@@
This plot allows for hovering over or clicking on the vertical bars (or on the line) in order to see the number of publications and citations received in a particular year.
Two sample outputs, produced by the MATLAB toolbox described below in this article, are shown in Fig. 1 (researcher R1) and Fig. 2 (researcher R2), and in both cases the researchers are mathematicians who have been listed in the Highly Cited Researchers lists in the past; their career lengths are similar.
@@
However, from 2011 on, the number of publications of R2 started to grow rapidly, giving the total of 303 publications in 2011–2023, reaching the peak of 45 papers in 2022, and the number of citations started to grow synchronously with the number of publications. The total number of co-authors of R2 during 2011–2023 was 148, from 35 countries.
The described differences between researchers R1 and R2 are obvious. It should be mentioned that the publication pattern of R1 also reflects the generally known observation that significant results are achieved in the first half of a mathematician’s career, and that the publication activity normally decreases with age. In the case of the researcher R2, one can observe the aforementioned snowball effect typical for the papermilling behavior.
Obviously, in the case of researcher R2, authoring 45 papers in mathematics in one year is not typical, and the synchronous growth of citation counts alongside the growth of paper counts is also atypical. On the contrary, in the case of researcher R1, the continuous growth of citations counts despite the decrease in the number of papers is normal and shows the long-lasting impact of R1’s works published in the first half of the considered time period.
The following features in patterns of publications and citations counts are possible indicators of the papermilling behavior:
@@
  • starting from some particular year, the number of publications starts to grow rapidly (usually monotonically), and reaches such numbers of papers per year that are unrealistic for a conscientious and responsible scientist;
@@
Besides these symptomatic signs of the overall publishing behavior, there are also symptoms that require scrutiny of each particular publication.
@@ 6 Basic quantitative indicators of the papermilling behavior
The number of publications in each year and the number of citations in the same year are two time series, and we can investigate their properties and mutual relationship. To quantify the previous observations, the following indicators can be used when examining the publication activity of researchers:
  • correlation between the publications counts and the citations counts; high value of the correlation coefficient indicates possible papermilling;
  • the delay between the citations counts and the publications counts; if the correlation is high and the delay is small, then this indicates possible papermilling;
  • unusually large numbers of publications in one year, with increasing trend, indicate possible papermilling;
  • low integrity index (see the next Section).
@@ 7 Integrity index
The main goal of a conscientious scientist is to discover new knowledge, and not to increase the number of his/her publications. In this respect, it is worth recalling Theodore von Kármán’s approach to comparing scientists [16, page 4]:
“If you define a great scientist as a man with great ideas, then you will have to rate Einstein first – he had four great ideas. In the history of science perhaps only Sir Isaac Newton is ahead of Einstein, because he had five or six ideas. All the other major scientists of our age are associated with just one, or at the most two great ideas.”
The growing numbers of published papers make it increasingly difficult to distinguish the valuable, novel, and original material from the materials that just increase the global number of publications.
The motivation for introducing the h-index by Jorge Hirsch [7] was to focus on the publications which received sufficient attention of the related scientific community reflected in citations, and not to focus on large numbers of publications that did not draw notable attention. However, the papermills allow a fast increase of the h-index by generating a large number of worthless papers and citations to such papers, also in worthless articles; this is the aforementioned snowball effect.
The solution can be introducing a quantitative indicator, which would implement von Kármán’s idea by lowering the motivation for increasing the number of publications by penalizing the unnecessary increase of the total number of publications.
As such indicator, I suggest the ratio of the h-index and the total number of publications. This can be considered as a kind of a signal-to-noise ratio, called the integrity index, and denoted as “$I$-index”.
If one author with h-index $H=32$ has 64 papers indexed in the Web of Science and the other author with the same h-index $H=32$ has 320 papers, then the integrity index of the first author is $I_{1}=32/64=0.5$ and the integrity index of the second author is $I_{1}=32/320=0.1$. This means the higher probability that the first author focuses on publishing novel original results and not on just publishing more papers.
In the example discussed above, the integrity index of researcher R1 is $I_{1}=86/196=0.44$, while the integrity index of researcher R2 is $I_{2}=34/384=0.08$. The integrity index can be used to differentiate conscientious scientists from the possible papermillers. In the case of the papermilling behavior, the integrity index is very low compared to scientists with the similar values of h-index.
@@ 8 The MATLAB function
A function for MATLAB, “researcherprofilewos.m”, has been created ([11]) to simplify analysis of the data obtained from the Web of Science. All figures in this article are produced with the help of this tool.
@@
  • maximal number of papers in one year;
@@
  • II-index (integrity index);
@@
  • total citations count;
@@
  • minimal number of papers in one year;
  • average number of papers in one year;
  • average number of citations per paper.
@@
The most important output is the coefficient of linear correlation between publications and citations. Strong correlation means that the citation dynamics is strongly linked to the publication dynamics. In the case of researcher R2 the coefficient of linear correlation between publications and citations is $r_{2}=0.94$, and there is no delay between the growth of the number of publications and the number of citations. For researcher R1, the correlation coefficient is the negative number $r_{1}=-0.23$, which means that R1’s citation counts grow despite decreasing publications counts.
9 A case study
As a case study, let us take a look at the list of 89 researchers identified by Clarivate as highly cited researchers in the field of mathematics and considered by Agricola et al. [2].
Unexpectedly, not all of them could be uniquely identified in the Web of Science by their Web of Science author identifiers [^2] This is strange, since inclusion of a researcher in the HCR list is considered as a distinction and must be always verifiable. This is especially important when there are numerous profiles with the same or very similar full names. Also, some of the researchers, who are listed in HCR 2019 in the field of mathematics, do not have such a mark in their profiles in the Web of Science.. The data for the remaining 82 researchers were processed using the aforementioned MATLAB function, and the results are presented in Figs. 36.
Figure 3: Integrity index versus correlation between the time series of papers and citations
Figure 4: Integrity, correlation between papers and citations, and maximum number of papers in one year.
Figure 5: Integrity index versus total number of papers, along with the power-law fit.
Figure 6: Maximum number of papers in one year versus total number of papers, along with the linear fit.
One of the main reasons for the papermilling behavior is the increase of the h-index, and the increase of the number of papers and the number of citations for this purpose results in strong correlation between the time series of publications and the time series of citations, practically without noticeable delay. As a result of inflating the total number of publications, the integrity index decreases. This is shown in Fig. 3.
The shaded rectangle corresponds to correlation greater than 0.5 and integrity index less than 0.3, and 68% of the 82 data points belong to this region. For the points in this region, the average total number of publications is around 487, the average maximum number of papers published in one year is around 52, and the average number of paper in one year is around 16.
For the data points outside the gray rectangle, the average total number of publications is around 159 (three times less), the average maximum number of papers published in one year is around 17 (also three times less), and the average number of paper in one year is around 6 (2.7 times less).
There are only three data points with high publications/citations correlation and integrity index greater than 0.3, but the corresponding total paper counts are around 100, and the maximum number of papers in one year is around 12.
In Fig. 4, the the maximum numbers of papers in one year are shown for the points from Fig. 3. This shows that the large numbers of papers in one year are typical for a large group of researchers with high correlation between their publications and citations time series.
From the Fig. 5, it is obvious that increasing the total number of publications means the decrease of the integrity index, and this relationship is very well fit by the power law. For the considered group of researchers, the equation of the fitting curve is $I(p)=1.7524\,p^{-0.5956}$, where $I(p)$ is the integrity index, and $p$ is the total number of papers.
The linear fit of the relationship between the total number of papers and the maximum number of papers in one year, shown in Fig. 6, confirms that the papermilling behavior leads to overproduction of publication units. For the considered group of researchers, the equation of the fitting line is m(p)=0.1151p3.3981m(p)=0.1151\,p-3.3981, where $m(p)$ is the maximum number of papers in one year, and $p$ is the total number of papers.
10 Reasons for the papermilling behavior
@@
There are various reasons for papermilling, and being listed in the lists of highly-cited researchers (HCR) is only one of them. The other reason is financial, because many universities link salaries and other financial benefits to the numbers of publications and citations, rankings, and also HCR listings. One of the most widespread reasons is pursuing promotion. Usually a candidate aspiring for promotion has to have a certain number of publications and citations in journals and conferences indexed in Web of Science or Scopus databases. The desire to achieve a promotion in a short time encourages some researchers to engage in papermilling activity and collaboration, because doing conscientious research and waiting for the eventual future citation impact delays the desired promotion.
@@
Two probable examples of such kind are shown in Fig. 7 (mathematics, Czech Republic; correlation between publications and citations is $r=0.9318$, maximum number of papers in one year is 28, total number of publications is 120, integrity index $I=0.15$) and Fig. 8 (mixed bag of environmental science, business, etc., Slovakia; correlation between publications and citations is $r=0.99691$, maximum number of papers in one year is 48, total number of publications is 142, integrity index $I=0.1479$).
@@
The damage done by the promotion-motivated papermilling should not be underestimated. This type of behavior is contagious and may strongly affect colleagues and students, pushing them to similar wrong and unethical practices. In addition, the promotion-motivated papermilling frequently helps such papermillers to get to influential positions or committees where they make important decisions about hiring, grants, and promotions of honest and conscientious researchers – and also of researchers similar to themselves.
Figure 7: Papermilling for promotion:
example 1 (mathematics, CZ).
Figure 8: Papermilling for promotion:
example 2 (environmental science, SK).
@@
11 Papermilling behavior
in other fields
The exclusion of the field of mathematics from the Clarivate’s lists of highly cited researchers does not mean that other fields are spared of papermilling, as the further two examples from Slovakia illustrate.
In Fig. 9 a well-documented example of a researcher, listed by Clarivate in the past as highly cited in the field of plant science, is shown. This case brings the picture similar to what we have seen in Fig. 2, Fig. 7, and Fig. 8, namely very low performance for long time in 1989–2013, and rapid growth after that: 11 papers in 2017, 14 in 2018, 19 in 2019, 33 in 2020, peaking at 90 papers in 2021, accompanied by the growth of citation counts from roughly 30 citations in 2013 to roughly 3100 citations in 2022. The number of co-authors is 1251, from 56 countries.
In Fig. 10, an example from the field of engineering is shown, very similar to the previous case, with very low performance until 2016, and extremely rapid growth after that: 3 papers and 23 citations in 2017, 11 papers and 56 citations in 2018, 32 papers and 307 citations in 2019, more than 13 papers and around 1300 citations in 2020, more than 110 papers and more than 2500 citations in 2021, 96 papers and more than 4300 citations in 2022. The number of co-authors is 1084, coming from 63 countries.
We can conclude that the papermilling behavior in other fields is also widespread and has the same signs.
12 What made the papermilling behavior possible?
It was the introduction of the APC (Article Processing Charges) and the payments for open access (OAF, open access fees), that radically changed the publication ecosystem and made the papermilling possible.
Before the introduction of the APC, the publishers were getting their revenue by selling subscriptions to their journals and reprints of the published papers. The authors provided their works to publishing houses for free or for a honorarium, which was a fraction of the revenue obtained by the publisher from selling the authors’ works. In such a relationship, the publishers were interested in publishing the top-notch papers, because one can hardly sell a meaningless or incorrect paper.
With the introduction of APC, a new source of the publisher’s revenue appeared: a huge number of researchers wanting to publish their articles for the price of APC.
The introduction of OAF made it possible for publishers to sign agreements with institutions and national authorities (such as ministries of education, science, research), in accordance with which those institutions or national authorities pay the OAF for the papers of their employees.
In the APC/OAF publishing model, publishers are – in the limit – interested in making all authors paying APC/OAF, because this will reduce to zero the risk of not selling any articles.
The arguments that the APC and OAF support open access publishing are very questionable, because the Arxiv.org and similar repositories also provide open access. Even more, such repositories definitely provide the so called “Diamond Open Access”, where the authors do not pay for publication and readers do not pay for reading.
It should be mentioned that although the publishers receive the APC and OAF money from the authors or the authors’ institutions, the reviewers of papers do their work for free – they are not paid for their work. This creates an incentive for manipulating the reviewing process, and even to creation of the “review mills”, as is documented in [9].